OpenStudy (anonymous):

five red bricks and two yellow bricks , 210 ,how many red bricks in the toy

OpenStudy (anonymous):

red=x+120 yellow=x 2(x+120)=5x 2x+240=5x x=80

OpenStudy (texaschic101):

what ?????